Most software manufactures offer a free help line that allows customers to call an 800 number to receive assistance with their problems. Because of the volume of calls, often customers are put on hold. Greeting Card Art, Inc. is proud of the fact that 93 percent of their calls for assistance are answered directly by a technician who can solve the customer's problem. To monitor this service claim, suppose Greeting Card Art selects a random sample of 25 recent calls.
A. Explain why this situation fits a binomial distribution.
B. What is the probability exactly 20 of the calls are answered directly by a technician?
C. What is the probability at least one of the calls is put on hold?

Answer by stanbon(75887)

Greeting Card Art, Inc. is proud of the fact that 93 percent of their calls for assistance are answered directly by a technician who can solve the customer's problem. To monitor this service claim, suppose Greeting Card Art selects a random sample of 25 recent calls.
A. Explain why this situation fits a binomial distribution.
Independent response on each call.
P(tech answer) = 0.93 for each call
n = 25
----------------------------------
B. What is the probability exactly 20 of the calls are answered directly by a technician?
P(x = 20) = 25C20(0.93)^20*(0.07)^5 = binompdf(25,0.93,20) = 0.0209
========================
C. What is the probability at least one of the calls is put on hold?
P(1<= x <=25) = 1 - P(x = none put on hold) = 1 - binompdf(25,0.07,0)
= 0.8370
=============
Cheers,
Stan H.
==============